OpenStudy (anonymous):

Box A has a volume of 12 cubic meters. Box B is similar to box A. To create box B, box A's dimensions were multiplied by five. What is the volume of box B?

OpenStudy (anonymous):

The first task you do is to put the box dimensions into a ratio (or fraction, whichever term you prefer). \[\frac{ A }{ B }\] Since we know we are multiplying our dimensions by five, we can write this equation: \[\frac{ 1 }{ 5 }\] To find the volume of the other box, we cube both sides of our ratio. \[\frac{ 1^{3} }{ 5^{3} }\] When you cube our ratio, you should get: \[\frac{ 1 }{ 125 }\] After figuring this out, we take the volume of Box A and multiply it by 125. \[12 \times 125 = 1,500cm ^{3}\]